40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 23 15:54:38 2010, kiwamu, Update, SUS, PRM and TT installed DSC_2275.JPG
    Reply  Fri Jul 23 18:55:25 2010, kiwamu, Update, SUS, PRM and TT installed 
Message ID: 3278     Entry time: Fri Jul 23 15:54:38 2010     Reply to this: 3281
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: PRM and TT installed 

[ Jenne, Koji and Kiwamu]

 We have installed the PRM and the tip-tilt (TT) in the BS chamber.

We have started the in-vac work which takes about a week.

Today's mission was dedicated to installing the PRM and two TTs, one for the PRC and the other for the SRC, on the BS table in the chamber.

The work has been smoothly performed and we succeeded in installation of the PRM and a TT for the PRC.

But unfortunately the other TT got broken during its transportation from Bob's clean room.

 


 (what we did)

 (1) opened the light door of the BS chamber.

 (2) moved the BS tower to the right position according to Koji's layout drawing.

    - Prior to this work we screwed down the earthquake stops so that the mirror is fixed to the tower. Also we disabled the watchdog.

    - When moving it we used an allen key as a lever with an screw as a fulcrum. This idea was suggested by Jenne and it really worked well.

     The reason why we used this technique is that if we slide the tower by hands the tower can't go smoothly and it may sometimes skips.

     After that we checked the postion from some reference screw holes by using a caliper and we made sure that it was on the right position.

 (3) removed all of the square-shaped mirrors.

    - After this removal the mirrors were wrapped by aluminum foils and put in a usual clear box.

 (4) removed some optics because they had made the chamber space crowded.

    - These were also wrapped by aluminum foils and put in the box. Later we will put them back to the BS table.

 (5) brought the PRM tower from the Bob's clean room  and put it on the BS table.

   - The position of the PRM were coarsely aligned since we still don't have any 1064 beam going through the PRM.

 (6) brought two TTs also from Bob's clean room and put one of the TTs on the table.  

   - The position of the installed TT was coarsely adjusted. 

   - After we brought them we removed the aluminum foils covering the TTs and we found the wire of a TT got broken.

     It may have been damaged during its transportation from Bob's room because it was fine before the transportation.

 (7) closed the door

 

(the next things to do)

  * Installation of the OSEMs to the PRM

  * Installation of the pick off mirror and its associated optics

  * Arrangement of  the pzt mirror

 

Attachment 1: DSC_2275.JPG  2.323 MB  | Hide | Hide all
DSC_2275.JPG
ELOG V3.1.3-